# Understanding Large Language Models in Your Pockets: Performance Study on COTS Mobile Devices

### v3

@@ IV Performance: Users’ Perspectives
In this section, we present performance metrics that affect user experience, including token throughput, memory footprint, and energy consumption. We specifically compare the key performance metrics across different types of processors, including CPUs, GPUs, and specialized AI accelerators (NPUs).
@@ IV-A1 Performance on CPUs
Figure 2: Inference Performance (llama.cpp)
Figure 3: Inference Performance (MLC LLM).
Figure 3 shows the performance of CPUs, including prefill speed and decoding speed with llama.cpp. Overall, there has been a consistent improvement in both prefill and decoding speeds with new-generation processors. The Apple M1, as a laptop-tier SoC, demonstrates the highest performance. The Dimensity $9300$, which ranks first among non-Apple SoCs, shows about a 1010% performance gap compared with the Apple M1 in the decode stage, but delivers only 6565% of the Apple M1’s throughput in the prefill stage. The Snapdragon 8 Gen $3$ (Xiaomi $14$ Pro) ranks third among all tested devices, showing 8080% throughput of the Dimensity $9300$. Comparing Snapdragon SoCs across different tiers, each successive generation shows approximately a $50\%$ improvement in prefill speed, while decode speed improves even more significantly, ranging from $80\%$ to $110\%$. We note that Snapdragon 870 (Huawei Matepad11 Pro) and Kirin 9000E (Huawei Matepad12.6 Pro) consist of the same cores, and the performance gap may result from higher CPU frequency on Kirin 9000E. Notably, a significant performance gap persists between Kirin SoCs and those from Snapdragon and MediaTek in terms of supporting LLMs.
Our results highlight a notable performance disparity in LLM inference across mobile CPUs. Specifically, top-tier SoCs like the Apple M1 or the Dimensity $9300$ demonstrate significantly better LLM support compared to mid-tier SoCs. This suggests that for optimal LLM performance, developers should focus on leveraging the most advanced processors available. Furthermore, top-tier SoCs support advanced machine instructions that can significantly boost performance. We will examine this potential in detail in Section V-A2.
We further evaluate the performance with two scenarios: one involving a prompt of $64$ tokens with a generation of $128$ tokens, and the other involving a prompt of $512$ tokens with a generation of $256$ tokens. The larger input matrix in the latter case implies greater computational intensity and memory resource demands. Figure 3 illustrates that inference speed decreases as the token length increases. The results may be attributed to the impact of DVFS. With longer prompts and generation tasks, the increased load causes the CPU frequency to throttle more aggressively. Across all tested devices, the performance degradation for both prefill and decoding is relatively consistent, ranging between $10\%$ and $20\%$. An exception is that the prefill speed achieves a $2.5\times$ improvement on the Apple M1, suggesting that matrix parallel computation on CPUs still has room for further optimization.
@@
In addition to latency, memory footprint is a also crucial metric. On all six devices, memory usage remains consistently around 3.83.8GB and shows minimal variation during inference once the model weights are loaded into main memory from external storage. This consistency is due to the fact that llama.cpp applies for a fixed-size memory block during initialization, which remains constant once allocation is complete. The allocated memory is then segmented into various components for different variables, including model weights, KV-cache, buffers, context and etc.
@@
LLMs running on mobile CPUs exhibit only slightly additional memory requirements beyond the model weights. Specifically, for a $4$-bit quantized lightweight LLM with 77B parameters, the maximum memory usage is no more than 44GB. Moreover, the parameters of many popular lightweight LLMs range from 11B to 55B, which further reduces memory requirement. Thus, as long as a mobile device has more than 44GB available RAM, it is sufficient to run these models locally.
@@ IV-A2 Performance on GPUs
Unlike CPUs, which are all from Arm, GPUs across vendors may have different architectures. Thus, it is challenging for developers to optimize for different GPUs. Figure 3 illustrates the end-to-end latency of MLC LLM across five high-tier GPUs from Arm Mali, Qualcomm Snapdragon, and Apple. It should be noted that Huawei MatePad 11 Pro and Huawei Nova 7 exhibit frequent crashes when running MLC LLM, which prevents the collection of reliable measurement results. Consequently, these two devices are excluded from the comparison. Among the tested GPUs, Mali GPUs, particularly the Mali-G78, exhibit significantly slow prefill speeds. Consequently, we only consider the performance with 64-token prompts. For the Mali-G78, due to its extremely poor prefill performance, we only present results for 32-token prompts as a reference.
Although the decode speed of the Apple M1 is comparable to that of the Adreno 750, its prefill speed is nearly $7\times$ higher. This result highlights the huge performance gap between Apple SoCs and non-Apple SoCs. The Adreno GPU consistently outperforms the Mali GPU in overall performance. Notably, Mali GPUs show nearly unusable prefill performance. When comparing products from the similar tier, the Adreno $750$ delivers $1.4\times$ faster decoding speed than the Mali-G720720. Similarly, the Adreno $730$ achieves $1.5\times$ faster decoding speed compared to the Mali-G7878. Although Mali GPUs have poor prefill performance, Mali-G720720 still achieves $1.5\times$ speedup in decoding than older-generation Adreno $730$. To look into this phenomenon, we will provide a detailed analysis in Section V-B. Memory usage across the GPUs is nearly identical at around 4.24.2GB, with a slight increase to 4.44.4GB observed on the Vivo Pad3 Pro.
@@
For non-Apple GPUs, despite their theoretical advantages for large language models (LLMs) due to their parallel floating-point computing capabilities, do not exhibit a clear advantage in practice. Counterintuitively, MLC LLM on GPUs performs worse in prefill speed compared to Llama.cpp on CPUs. The reason could be that optimizing LLMs on mobile GPUs is more complex than on CPUs due to the diversity of GPU architectures.
IV-A3 Performance on Specialized AI Accelerators
Specialized AI accelerators have recently been integrated into mobile SoCs, such as the Snapdragon Hexagon NPU, Kirin NPU, and MediaTek APU. However, these accelerators face several challenges, including limited support for dynamic shapes and floating-point operations, which complicates use for LLM inference [35]. The Hexagon NPU, being the only mobile NPU with an open instruction set architecture, has garnered significant research interest. Successful deployments of mobile LLM models on the Snapdragon 8 Gen 3 have been reported. We collect the throughput reported in [38](mllm) and PowerInfer-$2$ [35], and compare with the performance on CPU and GPU.
MLLM and PowerInfer-$2$ achieve roughly an order-of-magnitude improvement in prefill speed over CPU, with MLLM reaching $106$ tokens/s for a $64$-token prompt (versus $10.9$ tokens/s on CPU) and PowerInfer-$2$ achieving $88$ tokens/s for a $512$-token prompt (versus $8$ tokens/s on CPU). Given the memory-bound nature of the decode stage, a hybrid inference scheme—using the NPU for prefill and the CPU or GPU for decode—may be a suitable approach for future optimization.
Implication 3
Although manufacturers provide limited support for NPUs, early trials with the Hexagon NPU have shown a substantial advantage in inference tasks. In order to drive this forward, more efforts should be devoted to adapting models and algorithms to specialized AI accelerators.
TABLE II: Energy Consumption
Device Power Drain Prefill Decode
(Prompt, Generation) = (64, 128)Tokens (mAh/round) (token/s) (token/s)
Vivo Pad3 Pro $4.54$ $10.63$ $8.22$
Huawei Matepad12.6 Pro $8.28$ $6.67$ $4.34$
IV-B Battery Consumption
While the end-to-end latency of LLM inference on mobile CPUs shows promising potential, power consumption is also a key factor in determining the feasibility of these models for local, always-on user access.
Table II compares the power consumption and inference speed between the Vivo Pad3 Pro and Huawei MatePad 12.6 Pro. To obtain the power consumption, we read the data provided by the OS settings, which record and display the power consumption of a certain application. Old-generation devices are excluded from this comparison due to significant differences in battery capacity and power-saving policies compared to new-generation hardware. We conduct $20$ consecutive rounds of inference on the CPU using $6$ threads to balance performance and energy consumption, with a prefill length of $64$ and a generation length of $128$. Compared with the Kirin 90009000E (Huawei Matepad12.6 Pro), the Dimensity 9300 (Vivo Pad3 Pro) exhibits a $1.6\times$ speedup in prefill and a $1.9\times$ speedup in decoding, while consuming only $55\%$ of the power. These results suggest that top-tier SoCs are capable of delivering substantial energy savings alongside improved performance.
In recent years, most smartphones have been equipped with battery capacities ranging from 40004000mAh to 60006000mAh. At a power consumption rate of 99mAh per inference round, a device could theoretically support over $400$ rounds of inference. This result aligns with the findings in [8], reinforcing the feasibility of LLM inference on mobile devices. However, a significant rise in device temperature during inference suggests that LLM inference is a power-intensive task. We recorded an increase from 42.642.6°C to 66.866.8°C during a single inference on Huawei Matepad12.6 Pro. Thus, optimizing the power efficiency of LLMs on mobile devices is crucial for sustainable and long-term deployment.
(a) Throughput in Prefill Stage
(b) Throughput in Decode Stage
IV-C Parallel with Background Tasks (non-AI tasks)
@@
IV-D Parallel with AI Tasks
In addition to small workloads, there are scenarios where LLM inference must be executed alongside tasks that heavily consume compute or memory resources. For example, some image processing neural network models are deployed on mobile devices. It is important to evaluate whether the performance remains usable when these tasks run concurrently.
Figure 5 shows that when executed concurrently with other AI tasks, inference performance declines progressively as the number of parallel threads in the LLM inference process increases. Without additional parallel tasks, the best performance is achieved at $6$ threads, where CPU utilization reaches approximately $680\%$. However, when the YOLO-based application is running (introducing an additional $\sim 300\%$ CPU utilization), the optimal configuration shifts to $4$ threads, which yields the smallest relative degradation. Moreover, performance degradation in the decode stage becomes significantly more pronounced than in the prefill stage. The most critical case occurs at $8$ threads: prefill throughput drops by nearly $60\%$, while decode throughput collapses by more than $90\%$. In practice, this indicates that during the decode stage, inference with $8$ threads (equal to the number of CPU cores) becomes essentially unusable when YOLOv1111n runs concurrently.
Implication 4
With YOLOv11n consuming over $300\%$ of CPU resources, the highest inference performance is observed with fewer parallel threads. This suggests that the optimal thread count for practical scenarios is not fixed but instead depends on real-time CPU utilization. For AI workloads that demand substantial CPU capacity, parallel execution with LLMs presents significant challenges.
@@ V-A1 Multi-threads on CPU Cores
Most popular mobile SoCs utilize the “big.LITTLE” architecture for their CPUs, which balances performance with power efficiency. This configuration typically includes multiple cores organized into two distinct clusters: “big” (prime and performance cores) and “little” (efficiency cores), as illustrated in Table 3. The detailed CPU Specifications of Apple M1 are not publicly disclosed; it is known to comprise 4 performance cores and 4 efficiency cores. While it is commonly assumed that high-load tasks are best handled by the “big” cores for optimal performance, our tests reveal that the ideal core configuration can vary across the two stages of LLM inference.
Figure 6: Throughput with multi-threads.
@@
Figure 6 illustrates how these core configurations impact inference speed across the four tested devices. During the prefill stage, prompt processing speed is primarily driven by the big cores. The contribution of the little cores to acceleration depends on the performance gap between the big and little cores. On devices with limited computational power, additional efficiency cores can help accelerate inference. For instance, utilizing all cores delivers the best performance on the Kirin 9000E. However, for devices with more powerful big cores, adding little cores can actually reduce performance. For example, on the Snapdragon 8 Gen 3, incorporating two little cores results in a slowdown in inference speed. This highlights the importance of optimizing core configurations based on each device’s specific CPU capabilities to maximize efficiency. Interestingly, the Dimensity 9300 benefits from using all cores because of its All-Big-Core architecture [41]. On the Apple M1, we observe that even a single performance core achieves a prefill speed comparable to that of $8$ threads. Increasing the thread count does not improve CPU utilization; instead, it may introduce excessive context switching, where threads are frequently interrupted, leaving substantial idle intervals on CPUs and reducing CPU utilization.

### v4

@@ IV Performance: Users’ Perspectives
In this section, we present performance metrics that affect user experience, including token throughput, memory footprint, and energy consumption. We specifically compare the key performance metrics across CPUs and GPUs.
@@ IV-A1 Performance on CPUs
Figure 2: Inference performance on CPU (llama.cpp)
Figure 2 shows the performance of CPUs on different mobile devices, including prefill speed and decoding speed with llama.cpp. From the device perspective, there has been a consistent improvement in both prefill and decoding speeds with new-generation processors. Among the tested devices, the Apple M11 SoC exhibits the most superior performance as a laptop-grade processor. The Dimensity $9300$, while leading the non-Apple SoC category, trails the Apple M11 by approximately 1010% to 2020% during the decode stage. However, its exceptional performance significantly diminishes in the prefill stage, where it achieves 6060% of the Apple M11’s throughput for 64-token prompts and only 2020% for $512$-token prompts. The Snapdragon $8$ Gen $3$ follows as the third-best device, delivering roughly 8080% of the throughput observed in the Dimensity $9300$. Additionally, a significant performance gap exists between Kirin SoCs and their Snapdragon and MediaTek counterparts in terms of LLM support. From the model perspective, we draw the following observations:
Linear Scaling of Computational Overhead: The computational cost exhibits a nearly linear relationship with increases in model size. Across different devices and configurations, the throughput ratios closely mirror the model size ratios within the same family. For instance, after applying $4$-bit quantization, the Qwen33-44B model, which is roughly twice the size of Qwen33-11B, demonstrates a prefill speed that is only $0.3\times$ to $0.5\times$ of the Qwen33-11B when tested on the same device and under the same configuration.
Stage-Specific Sensitivity: The growth of model parameters has a more noticeable effect on prefill performance than on decoding, as the prefill stage is compute-bound. Larger models, with increased hidden dimensions and depth, result in greater computational load. For instance, with Llama3 models, the prefill performance decreases by 55% to 1010% more than decoding when scaling from 11B to 33B. This trend persists across devices, with larger models consistently exhibiting a similar pattern of performance degradation compared to their smaller counterparts. This pattern also holds for Qwen33 models.
Architectural Influence on Throughput: Model architecture plays a critical role in determining both parameter distribution and execution efficiency. This is particularly evident when comparing models with nearly identical parameter size, such as Llama3.23.2-11B and Gemma33-11B. Despite both models occupying approximately 760760MB post-quantization, their performance metrics diverge across different inference stages. Specifically, Llama3.23.2-11B demonstrates slower prefill speeds but superior decoding performance compared to Gemma33-11B. These differences underscore the inherent architectural trade-offs between these two distinct inference phases.
Sequence Length and Hardware Elasticity: Increasing the input sequence length intensifies both computational demand and memory pressure. On most non-Apple devices, extending the prompt length from 64 to 512 tokens results in a performance degradation exceeding $10\%$. In contrast, the Apple M1 architecture demonstrates remarkable elasticity: while decoding speed remains stable, prefill speed improves by a factor of $2.5$ to $3$ with longer prompts, highlighting significant potential for handling complex computations.
Figure 3: Inference performance (MLC LLM) on GPU with 64-token prompt
@@
In addition, memory footprint is also a crucial metric. The memory usage closely corresponds to the size of the model weights. On all six devices, memory usage remains consistently around 3.83.8GB when running Llama22-77B and below 11GB when running Gemma33-11B. Additionally, it shows minimal variation during inference once the model weights are loaded into main memory from external storage. This consistency is due to the fact that llama.cpp applies for a fixed-size memory block during initialization, which remains constant once allocation is complete. The allocated memory is then segmented into various components for different variables, including model weights, KV-cache, buffers, context and etc.
@@
LLMs running on mobile CPUs require only slightly additional memory beyond the model weights. Specifically, for a $4$-bit quantized lightweight LLM with 77B parameters, the maximum memory usage is no more than 44GB. Moreover, the parameters of many popular lightweight LLMs range from 11B to 55B, which further reduces memory requirement. Thus, as long as a mobile device has more than 44GB available RAM, it is sufficient to run these models locally.
@@ IV-A2 Performance on GPUs
Unlike CPUs, which are all from Arm, GPUs across vendors may have different architectures. Thus, it is challenging for developers to optimize for different GPUs. Figure 3 illustrates the end-to-end latency of MLC LLM across five high-tier GPUs from Arm Mali, Qualcomm Snapdragon, and Apple. It should be noted that Huawei MatePad 11 Pro and Huawei Nova 7 exhibit frequent crashes when running MLC LLM, which prevents the collection of reliable measurement results. Consequently, these two devices are excluded from the comparison. Among the tested GPUs, Mali GPUs, particularly the Mali-G78, exhibit significantly slow prefill speeds. Consequently, for ease of comparison, we focus only on the performance with 64-token prompts. It should be noted that the test for Llama2-7B on the Mali-G78 was conducted using 32-token prompts due to crashes that frequently occurred with longer prompts.
Hardware Perspective: The results highlight a significant performance gap between Apple SoCs and non-Apple SoCs. For models around 11B, the throughput on the Apple M1 is approximately $2\times$ faster than on the Adreno $750$. As for non-Apple SoCs, the Adreno GPU consistently outperforms the Mali GPU in overall performance across different models. Despite Mali GPUs showing poor prefill performance, the Mali-G720720 achieves a slight speedup in decoding compared to the older-generation Snapdragon SoC, i.e., the Adreno $730$.
For the two Mali GPUs, both prefill and decode speeds are slower than on the CPUs. The slowdown is especially significant in the prefill stage. For models larger than 44B, the prefill speed drops below $5$ tokens/s on the Mali-G720720, despite it having the highest theoretical computational capability among all non-Apple devices (see Table IV). This phenomenon and its underlying causes will be further explored in Section V-B.
Model Perspective: From the model perspective, we observe that for both Apple and Snapdragon GPUs, the prefill speed improves several times over CPU on the same device. For instance, on the Apple M11, the prefill speed of all models is $3$ to $4\times$ faster. However, the decode speed on the GPU is quite similar to that on the CPU, indicating that while the GPU offers a clear advantage in computation and can greatly accelerate prefill, the decode stage remains the bottleneck that limits overall throughput.
In the prefill stage, the Llama33 models show the highest improvement on both Apple and Snapdragon devices. For example, on the Snapdragon 88+ Gen11, Llama3.23.2-33B achieves a $4\times$ improvement in prefill speed. For the 11B models from Gemma33 and Llama3.23.2, despite the similar model sizes, the prefill acceleration for Llama3.23.2-11B is higher than Gemma33-11B on the same device. This suggests that the architecture of Llama3.23.2 is more efficient for parallel computing.
Additionally, memory usage across the GPUs is nearly identical, around 4.24.2GB when running Llama22-77B, with a slight increase to 4.44.4GB observed on the Vivo Pad33 Pro. However, as the prompt length and prefill chunk size increase, memory usage grows accordingly, showing higher memory demand compared to inference on CPU.
@@
Although GPUs do not exhibit a clear advantage for large models (Llama22-77B), they show a significant improvement in prefill for smaller models (e.g., 11B, 33B). It is clear that hybrid execution, such as performing prefill with the GPU and decoding with the CPU, is an effective strategy. Furthermore, from the suboptimal performance observed on Mali GPUs, GPU acceleration often requires device-specific tuning to overcome architectural disparities and achieve competitive efficiency.
IV-B Parallel with Background Tasks (non-AI tasks)
@@
IV-C Parallel with AI Tasks
In addition to small workloads, there are scenarios where LLM inference must be executed alongside tasks that heavily consume compute or memory resources. For example, some image processing neural network models are deployed on mobile devices. It is important to evaluate whether the performance remains acceptable when these tasks run concurrently.
Figure 5 shows that when executed concurrently with other AI tasks, inference performance declines progressively as the number of parallel threads in the LLM inference process increases. Without additional parallel tasks, the best performance is achieved at $6$ threads, where CPU utilization reaches approximately $680\%$. However, when the YOLO-based application is running (introducing an additional $\sim 300\%$ CPU utilization), the optimal configuration shifts to $4$ threads, which yields the smallest relative degradation. Moreover, performance degradation in the decode stage becomes significantly more pronounced than in the prefill stage. The most critical case occurs at $8$ threads: prefill throughput drops by nearly $60\%$, while decode throughput collapses by more than $90\%$. In practice, this indicates that during the decode stage, inference with $8$ threads (equal to the number of CPU cores) becomes essentially unacceptable when YOLOv1111n runs concurrently.
Implication 3
With YOLOv11n consuming over $300\%$ of CPU resources, the highest LLM inference performance is observed with fewer parallel threads. This suggests that the optimal thread count for practical scenarios is not fixed but instead depends on real-time CPU utilization. For AI workloads that demand substantial CPU capacity, parallel execution with LLMs presents significant challenges.
Implication 4
(a) Throughput in Prefill Stage
(b) Throughput in Decode Stage
@@ IV-C Parallel with AI Tasks
IV-D Accuracy vs Latency
We evaluate six pretrained models without fine-tuning across three benchmarks to assess various inference capabilities, including the MMLU [27] and ARC [15] for general problem-solving, and TruthfulQA [38] for reliability and safety. As shown in Figure 6, within a given model family, performance correlates positively with parameter scale. However, architectural differences among families introduce significant performance disparities. For instance, Qwen33 consistently outperforms the Llama and Gemma33 series. Notably, Qwen33-1.71.7B shows a substantial advantage over Llama22-77B, surpassing its MMLU accuracy by nearly 1515% despite the latter having a larger parameter count. While Llama3.23.2-11B and Gemma33-11B demonstrate nearly identical accuracy in line with their similar scales, they are substantially outperformed by Qwen33-1.71.7B.
Regarding quantization schemes, most models maintain performance comparable to the F1616 baseline across various bit-widths, with significant degradation occurring only at extremely low levels such as $3$-bit. Interestingly, transitioning from Q33_K_M to Q44_K_M on TruthfulQA yields divergent outcomes: while the accuracy of Qwen33-1.71.7B improves, Llama22-77B and Gemma33-11B experience declines exceeding 11%. These findings suggest that higher bit-widths do not necessarily enhance performance, underscoring the necessity of post-quantization evaluation to identify the optimal configuration for specific tasks.
Figure 6: Model accuracy on different tasks (TruthfulQA, MMLU, ARC)
Figure 7: Inference throughput for different quantization types on CPU (llama.cpp)
From the efficiency perspective, as shown in Figure 7, $4$-bit quantization strikes the optimal balance by achieving peak or near-peak speeds in both prefill and decoding stages while maintaining robust inference accuracy. Leveraging the advanced instruction optimizations (see Section V.A), Q44_0 achieves a significantly higher throughput on the Xiaomi $14$ Pro compared to other quantization schemes. For small-scale LLMs, such as Gemma33-11B, 88-bit quantization serves as a viable alternative to enhance response quality without sacrificing throughput. In the compute-bound prefill stage, both $4$-bit and $8$-bit schemes exhibit clear advantages as their bit-widths align with the standard 8-bit byte structure. Schemes that do not align with this byte structure necessitate additional bitwise operations for dequantization, thereby increasing computational overhead. Conversely, in the memory-bound decoding stage, despite the smaller memory footprint of 3-bit quantization, 4-bit quantization generally offers slightly higher throughput than its 3-bit alternative.
@@ IV-D Accuracy vs Latency
The architectural diversity of mobile hardware necessitates a hardware-aware deployment strategy rather than a universal quantization standard. For top-tier devices with instruction-level acceleration, the Q4_0 scheme offers an optimal balance by maximizing execution speed while preserving accuracy. In contrast, for resource-constrained platforms, pairing a smaller model with 8-bit quantization provides a superior compromise between response quality and practical throughput. Ultimately, on-device benchmarking of quantization is crucial to align the model with specific hardware constraints and the requirements of the target task, optimizing overall performance.
@@ V-A1 Multi-threads on CPU Cores
Most popular mobile SoCs utilize the “big.LITTLE” architecture for their CPUs, which balances performance with power efficiency. This configuration typically includes multiple cores organized into two distinct clusters: “big” (prime and performance cores) and “little” (efficiency cores), as illustrated in Table III. The detailed CPU Specifications of Apple M1 are not publicly disclosed; it is known to comprise 4 performance cores and 4 efficiency cores. While it is commonly assumed that high-load tasks are best handled by the “big” cores for optimal performance, our tests reveal that the ideal core configuration varies across the two stages of LLM inference.
Figure 8: Throughput with multi-threads (llama.cpp with Llama22-77B)
(a) Xiaomi14 Pro (Snapdragon 8 Gen3)
(b) Vivo Pad3 Pro (Dimensity 9300)
Figure 9: Inference Performance with INT8 Matrix Multiplication (llama.cpp)
@@
Figure 8 illustrates how these core configurations impact inference speed across the four tested devices. During the prefill stage, prompt processing speed is primarily driven by the big cores. The contribution of the little cores to acceleration depends on the performance gap between the big and little cores. On devices with limited computational power, additional efficiency cores can help accelerate inference. For instance, utilizing all cores delivers the best performance on the Kirin 9000E. However, for devices with more powerful big cores, adding little cores can actually degrade performance. For example, on the Snapdragon 8 Gen 3, incorporating two little cores results in a slowdown in inference speed. This highlights the importance of optimizing core configurations based on each device’s specific CPU capabilities to maximize efficiency. Interestingly, the Dimensity 9300 benefits from using all cores because of its All-Big-Core architecture [41]. On the Apple M1, we observe that even a single performance core achieves a prefill speed comparable to that of $8$ threads. Increasing the thread count does not improve CPU utilization; instead, it may introduce excessive context switching, where threads are frequently interrupted, leaving substantial idle intervals on CPUs and reducing CPU utilization.